OpenStudy (anonymous):

i need help with polynomials, subtract the polynomials.... (r-1/12s) - (1/2r - 5/12s - 4/11)

hero (hero):

(r-1/12s) - (1/2r - 5/12s - 4/11) =r - 1/12s -1/2r + 5/12s + 4/11 = 1/2r + 4/12s + 4/11